Determine the producer surplus when equilibrium price is Pe=5. Given the supply function of p=S(q)= q + 2


1
Expert's answer
2021-09-21T15:05:06-0400

At equilibrium point

"5=q+2 \\

q =3"

The producer surplus is

"PS= pq - \int^q_0 s(q)dq \\

PS = 5 \times 3 -\int^3_0 (q+2)dq \\

PS = 15 -[\frac{q^2}{2} + 2q]^3_0 \\

PS = 15 -((4.5+6) -(0)) = 15 -10.5 = 4.5"